OpenStudy (anonymous):

(2/(sqroot5)-2(sqroot6))/(1+2/(sqroot5)*2/(sq6)

OpenStudy (anonymous):

Simplify

OpenStudy (anonymous):

\[\frac{ \frac{ 2 }{ \sqrt{5} } - 2\sqrt{6} }{1+ \frac{ 2 }{ \sqrt{5} } - \frac{ 2 }{ \sqrt{6} }}\]